The length of a rectangular garden is 7 ft less than 3 times its width. If the perimeter of the garden is 42ft, find the deminsions of the garden.
Could I please get some help with this. Thank you

Answer by checkley71(8403)

(3X-7)=THE LENGTH & X=WIDTH THEN 2(3X-7)+2X=42 OR 6X-14+2X=42 OR 8X=42+14 OR
8X=56 OR X=56/8 OR X=7 THIS IS THE WIDTH THEN THE LENGTH IS 3*7-7 OR 21-7=14
THEN 3*7-7+2*14=42 OR 21-7+28=42 OR 49-7=42 OR 42=42